Determine the margin of error for a confidence interval to estimate the population mean with n = 35 and ? = 49 for the following confidence levels. a) 91% b) 94% c) 97% a) with a 91% confidence level, the margin of error is (Round to two decimal places as needed.)
Answer #1

Margin of Error formula:

E = z\sigma/\sqrt{n}

a) For 91% confidence level, z = 1.695

Hence,

E = 1.695*49/\sqrt{35} = 14.04

b) For 94% confidence, z = 1.88

Hence,

E = 1.88*49/\sqrt{35} = 15.57

c) For 97% confidence, z = 2.17

Hence,

E = 2.17*49/\sqrt{35} = 17.97